OpenStudy (anonymous):

if 20 capacitors, each of 100 microFarad, were connected in parallel across a 12-V battery, what would be the total energy stored in the capacitors?

OpenStudy (marco26):

\[W=\frac{ 1 }{ 2 }CV ^{2}\]

OpenStudy (goformit100):

For Solving this Question, I would like to ask you that, do you have the detailed conception of : Capacitor and Capacitance , not just only the formula

OpenStudy (anonymous):

total capacitance:20x100x10^-6F=2x10^-3F v=12v E=1/2cv^2=1/2x2x10^-3x12=12x10^-3J=12mJoules.